# Cam-2-Cam: Exploring the Design Space of Dual-Camera Interactions for Smartphone-based Augmented Reality

### v2

@@ Abstract
Introduction
Designing dynamic interactions for smartphone-based AR has traditionally been constrained by single-camera configurations typical in mobile devices, which either solely rely on touch-based input (e.g.,Pokemon Go [32]) or only realize the front-facing selfie camera for AR content (e.g.,Snapchat facial filters [42]) [22, 11]. Other AR form factors such as HMDs have used multiple cameras to extend the interaction space, typically allowing for more of the users’ physical gestures to be captured and to expand their AR environment where users can have more immersive experiences, a common goal for AR interfaces. In contrast, smartphone AR usually only leverages either the front-facing or the rear-facing camera and not both in tandem, which constrains the AR environment for smartphone AR [36, 38, 26, 55]. For this work, we adopt the definition of immersion presented by Agrawal et al., which states that immersion is “A state of deep mental involvement in which the individual may experience disassociation from the awareness of the physical world due to a shift in their attentional state.” [2]. To use both cameras on the smartphone meaningfully, we delegate the front-facing camera to capture a touchless physical gesture since it already faces the user, and we render the virtual reaction in the rear camera feed, situating elements in front of the user along their line of sight. This approach builds on how humans would naturally interact with objects placed in front of us in a physical space [25, 11, 37, 36, 18]. We aim to investigate the interaction possibilities that can motivate research opportunities centered on circumventing the limited field
@@ Introduction
Figure 1: Cam-2-Cam leverages both rear-facing camera and front-facing camera, traditionally used independently, to extend the interaction space in smartphone-based AR applications (Left). We explore the design space through three unique applications–A.Face TriggAR (wink gesture), B. Mouth Craft (mouth-open gesture), and C. Mirror ThrowAR (free-hand manipulation gestures) (Right).
@@
of view (FOV) and small screen size, which can hinder the AR interaction experience on smartphones [37, 18, 55].
@@
To this end, we present the Cam-2-Cam interaction concept, implemented in three dual-camera AR applications that utilize different camera input modalities (face or hand gestures) and camera configurations (alternating or simultaneous capture) to capture the breadth of the dual-camera interaction space. Each application provides a unique example of interaction types feasible with dual-camera systems: Face TriggAR uses a wink gesture (i.e. one-eye-closed) to simulate aiming and shooting mechanics, Mouth Craft enables block placement with mouth gestures, and Mirror ThrowAR integrates free-hand throwing detection across front and rear cameras (see figure 1). In addition to expanding on gesture-based interactions, Cam-2-Cam draws on the concepts of responsive multimodal feedback, incorporating haptic, visual and auditory cues to enhance users’ perception of control and cohesion across dual-camera setups, informed by past work on multimodal feedback integration in AR systems [11, 36, 3, 4].
@@
Our primary goals of this study are to construct design lessons from participants’ responses that can connect the two physical spaces captured by the front and rear-facing cameras, creating a cohesive or unified AR interaction space. We define cohesive throughout this paper as a sense of connection between states or scenes, which is analogous to research centered on virtual transitions in XR and camera techniques in cinema [34, 27, 8, 39, 40, 45]. Due to the novelty of dual-camera setups for smartphone-based AR interactions, our study takes the form of an exploratory experiment where we use varying input modalities and camera techniques to cover the breadth of the space, highlight key design lessons in expanding the smartphone AR interaction space, and the significance of a larger interaction space concerning the goals of AR research.
@@
We release the three Cam-2-Cam applications as open source to support future design probes: [anonymized for submission].
1 Related Work
@@
1.1 Smartphone AR Interactions with Touchless Gestural Input
@@
Touchless gestural input has been widely explored in smartphone-based AR to provide a user experience that is natural to users based on familiar actions they’d perform in real life. For instance, Brasier et al. presented a study to offload parts of the on-screen GUI to allow users to interact with hand gestures, allowing more screen content to be visible [11]. Qian et al. expanded on this and designed a free-hand manipulation system to manipulate AR objects using the rear-facing camera on a smartphone where they found through an iterative design process that haptic, visual, and sound cues assisted users in their AR tasks [36]. Additionally, Loorak et al. presents hand and face interactions to interact with the on-screen UI, demonstrating users’ preference for this touchless interaction versus a touch-based interface when taking selfie photos [25]. The literature above illustrates the potential of touchless interactions, but they also share drawbacks they encountered when designing an AR experience for a single camera, as well as limited screen space with a narrow field of view.
1.2 Expanding the Interaction Space of Smartphones
@@
Expanding the camera view of smartphones has been explored to expand the user interaction space and environmental awareness. [18] introduce GlassHands, which uses reflections from sunglasses to enable interaction around unmodified mobile devices. Yeo et al. present OmniSense, a smartphone system enhanced with an omnidirectional camera to provide better around-the-body awareness [52]. These approaches expand the field of view of mobile devices, allowing for more immersive, spatially demanding interactions – these findings share common insights with literature centered on expanding the input space of head-mounted displays with multi-camera configurations [55, 44, 7, 40]. Cam-2-Cam similarly seeks to broaden the interaction capabilities of smartphones, but focuses on using the existing front and rear cameras to capture different types of input and output without additional hardware.
1.3 Dual-Camera Applications in Mobile Tracking and Smartphone AR
@@
Prior work has demonstrated how smartphone cameras working synchronously can enhance real-world conditions experiences by providing additional sensing and improving environmental understanding. For instance, Babic et al. introduced SIMO, a system that uses the front camera for face and head tracking and the rear camera to estimate depth from a distant display to approximate body tracking [9]. Nagai et al. presented HandyGaze, which uses gaze and head tracking via the front camera combined with the rear camera and pre-scanned 3D model of the environment to localize users’ gaze direction to identify artwork and present relevant information on a website in a mobile browser [30].
@@
Zhao et al. used both the front-facing and rear-facing cameras to capture more accurate ambient lighting, enabling more realistic shaders and textures for AR try-on applications [53]. To the best of our knowledge, this is the only prior work that explicitly employs a dual-camera configuration for augmented reality on smartphones. However, while Zhao et al. focus on visual fidelity, the role of dual-camera setups for smartphone AR interaction techniques remains largely unexplored. Our work addresses this gap by investigating the interaction design potential of dual-camera configurations, extending the interaction space by delegating specific functions to each camera, without modifying the smartphone’s form factor.
@@
2 Cam-2-Cam Design
Face TriggAR, Mouth Craft, and Mirror ThrowAR serve as design probes for the Cam-2-Cam interaction concept. Each application expands the smartphone AR input space by delegating the front-facing feed gestural input and the rear camera feed to situate AR content. We delegated the rear-facing camera to display the resulting AR elements from users’ gestural actions captured by the front-facing camera because AR research and mainstream AR applications are primarily concerned with interacting with the physical world in front of the user (i.e., along the line of sight of the user). Additionally, our implemented applications incorporate audio, visual, and haptic feedback to enhance user engagement and provide immediate multimodal confirmation of actions. This design approach aligns with prior mobile AR research [26, 36, 38, 50] emphasizing the benefits of multimodal feedback and touchless gestures in improving screen occlusion caused by limited screen real estate and enhancing user perceived control in smartphone-based AR applications.
2.1 Face Gesture Task Preferences
@@
Figure 2: Bar chart shows the number of participants selecting each gesture as their preferred dual-camera AR facial input after trying out projectile shooting and block object placement prototypes.
@@
2.2 Mobile Platform Considerations
@@
To enable hand gestures for free-hand manipulation of AR content we used the Portal-ble free-hand AR library due to its efficacy in manipulating AR objects as shown in several studies [36, 38, 26]. The Portal-ble library is only supported on Android, so we implemented Mirror ThrowAR application on Android (using the latest Google Pixel 9) rather than iOS.
@@
2.3 Face TriggAR
Face TriggAR is developed with Swift, using frameworks for AR mobile scenes (ARKit and RealityKit [6]), built for iOS smartphones and later deployed on the iPhone 15. It uses world, motion, and scene tracking with a single multi-camera capture session with simultaneous front and rear-cameras for a paintball target practice simulation.
2.3.1 Scene Recognition and Target Generation:
@@
Figure 3: Face TriggAR: (1) Displays a zoomed-in view of the user performing the wink gesture registered by the front-facing camera. (2) Shows the resulting projectiles (virtual paintballs) being shot along the user’s line of sight and visualized in the rear camera’s view.
2.3.2 Triggering Projectile Launches:
@@
2.4 Mouth Craft
Mouth Craft is an AR application designed for the iPhone 15 that uses ARKit and RealityKit frameworks with Swift to create a touchless interaction system driven by mouth gestures for object placement. Users can build virtual 3D structures by placing Minecraft-styled [28] grass blocks into their scene environment with the front-facing camera continuously tracking facial expressions. This design allows users to act in real-time without transition delays, making the AR environment feel more immediate and responsive. When a user opens their mouth, the blocks are situated in the AR scene and rendered to the rear camera’s feed.
2.4.1 Mouth Input:
@@
Figure 4: Mouth Craft: Starting from left to right, we demonstrate the (1) user performing the ‘mouth open’ gesture that is captured within the FOV of the front-facing camera and a detected plane is captured within the FOV of the rear-facing camera. (2) A block is placed on the detected plane and situated in the physical space captured by the rear-facing camera.
2.4.2 Block Scene Placement:
@@
2.5 Mirror ThrowAR
2.5.1 Throwing Implementation:
Mirror ThrowAR builds upon an open-source smartphone AR system, Portal-ble, to enable free-hand throwing detection [36]. The system uses the Google Pixel 9’s front-facing and rear-facing camera to track the user’s spatial hand positions and convert them to 3D coordinates in the AR application in real time. These coordinates are used to detect overhand throwing gestures for our experiment.
We consider a three-dimensional Cartesian system, where $x$ is the horizontal axis, $y$ is the vertical axis relative to the virtual target generated on the wall, and $z$-axis is perpendicular to both $x$ and $y$, pointing outwards away from the cameras making the $z$-axis the forward axis in respect to the target location. Given these three axes, we can define three perpendicular planes: the $x$-$y$ plane, which describes the wall where the target is anchored; the $y$-$z$ plane that defines the vertical space between the participant and the target; and the $x$-$z$ plane, which identifies the horizontal space between the participant and the wall (see equation 1). We used ballistic motion calculations adapted from [14] :
$\Theta_{i}=\arctan\frac{\left(S_{i}^{{}^{2}}+\sqrt{S_{i}^{4}-G\left(Gx^{2}+2S_{i}^{2}\right)}\right)}{G}$ (1)
@@ 2.5.1 Throwing Implementation:
Figure 5: Mirror ThrowAR users hold their hand out in a pinch or ball grabbing gesture to generate a Poké Ball fixed to their hand visible on their smartphone. Then, they throw the ball toward their smartphone into the rear camera AR scene to try to hit the Pikachu target.
@@
3 Formative Evaluation
3.1 Participants
@@
Our recruitment goal was 30 participants as this is often sufficient to reach saturation of themes in qualitative experiments, and a higher number typically results in more nuanced insights [19, 21]. The recruited participants included 12 females and 18 males between the ages of 18 and 60 years old ($\mu=26.97$, $\sigma=9.41$). Participants also provided their experience with smartphone AR during our semi-structured interview and rated on a 5-point Likert scale (1 = No Experience, 5 = Very Experienced). The resulting levels of experience can be summarized as ($\mu=1.93$, $\sigma=0.89$).
3.2 Study Protocol
@@
Each session began with experimenters demonstrating the applications to ensure participants were familiar with the setup and controls. In Face TriggAR, participants first scanned a QR code anchor image to generate a fixed AR target, spending the initial moments aiming at this anchored target. Afterwards, they could manually reset the target location such as on a table or coach, allowing them to explore aiming from various angles and even move behind the target to observe it from different perspectives. Using a wink gesture to aim and shoot at each target, participants had two minutes to interact with the scenes.
@@
3.3 Data Analysis
Experimenters conduct semi-structured interviews with the following qualitative questionnaire based on the AR Design Heuristics presented by Endsley et al. and tailored to fit our research questions displayed in table 1 [15]. The first and third authors carried out the thematic analysis by initially performing an open coding procedure followed by axial coding independently. Afterward, they reconciled differences in the final code groups and formed themes as seen in table 2.
Survey Questions AR Design Heuristics [15]

### v5

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025conference: Conditionally Accepted to ACM Symposium on Spatial User Interaction (SUI 2025); Nov. 10–11, 2025; Montreal, QC, CAccs: Human-centered computing User Experienceccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Gestural Input
@@
Figure 1. Cam-2-Cam leverages both rear-facing camera and front-facing camera, traditionally used independently, to extend the interaction space in smartphone-based AR applications (Left). We explore the design space through three unique applications–A.Face TriggAR (wink gesture), B. Mouth Craft (mouth-open gesture), and C. Mirror ThrowAR (free-hand manipulation gestures) (Right).
1. Introduction
Designing dynamic interactions for smartphone-based AR has traditionally been constrained by single-camera configurations typical in mobile devices, which either solely rely on touch-based input (e.g.,Pokemon Go (Niantic, 2024)) or only realize the front-facing selfie camera for AR content (e.g.,Snapchat facial filters (Snap, 2024)) (Kim and Lee, 2016; Brasier et al., 2021). Other AR form factors such as HMDs have used multiple cameras to extend the interaction space, typically allowing for more of the users’ physical gestures to be captured and to expand their AR environment where users can have more immersive experiences, a common goal for AR interfaces. In contrast, smartphone AR usually only leverages either the front-facing or the rear-facing camera and not both in tandem, which constrains the AR environment for smartphone AR (Qian et al., 2019, 2020b; Ma et al., 2023; Zhu and Grossman, 2020). For this work, we adopt the definition of immersion presented by Agrawal et al., which states that immersion is “A state of deep mental involvement in which the individual may experience disassociation from the awareness of the physical world due to a shift in their attentional state.” (Agrawal et al., 2019).
To use both cameras on the smartphone meaningfully, we delegate the front-facing camera to capture a touchless physical gesture since it already faces the user, and we render the virtual reaction in the rear camera feed, situating elements in front of the user along their line of sight. This approach builds on how humans would naturally interact with objects placed in front of us in a physical space (Loorak et al., 2019; Brasier et al., 2021; Qian et al., 2020a, 2019; Grubert et al., 2016). We aim to investigate the interaction possibilities that can motivate research opportunities centered on circumventing the limited field
@@ 1. Introduction
of view (FOV) and small screen size, which can hinder the AR interaction experience on smartphones (Qian et al., 2020a; Grubert et al., 2016; Zhu and Grossman, 2020).
@@
To this end, we present the Cam-2-Cam interaction concept, implemented in three dual-camera AR applications that utilize different camera input modalities (face or hand gestures) and camera configurations (alternating or simultaneous capture) to capture the breadth of the dual-camera interaction space. Each application provides a unique example of interaction types feasible with dual-camera systems: Face TriggAR uses a wink gesture (i.e. one-eye-closed) to simulate aiming and shooting mechanics, Mouth Craft enables block placement with mouth gestures, and Mirror ThrowAR integrates free-hand throwing detection across front and rear cameras (see figure 1). In addition to expanding on gesture-based interactions, Cam-2-Cam draws on the concepts of responsive multimodal feedback, incorporating haptic, visual and auditory cues to enhance users’ perception of control and cohesion across dual-camera setups, informed by past work on multimodal feedback integration in AR systems (Brasier et al., 2021; Qian et al., 2019; Alem et al., 2011; Aliprantis et al., 2019).
@@
Our primary goals of this study are to construct design lessons from participants’ responses that can connect the two physical spaces captured by the front and rear-facing cameras, creating a cohesive or unified AR interaction space. We define cohesive throughout this paper as a sense of connection between states or scenes, which is analogous to research centered on virtual transitions in XR and camera techniques in cinema (Pointecker et al., 2022; Mayer et al., 2024; Auda et al., 2023; Roo and Hachet, 2017; Sakashita et al., 2023; Tseng and Tseng, 2013). Due to the novelty of dual-camera setups for smartphone-based AR interactions, our study takes the form of an exploratory experiment where we use varying input modalities and camera techniques to cover the breadth of the space, highlight key design lessons in expanding the smartphone AR interaction space, and the significance of a larger interaction space concerning the goals of AR research.
@@
2. Related Work
@@
2.1. Smartphone AR Interactions with Touchless Gestural Input
@@
Touchless gestural input has been widely explored in smartphone-based AR to provide a user experience that is natural to users based on familiar actions they’d perform in real life. For instance, Brasier et al. presented a study to offload parts of the on-screen GUI to allow users to interact with hand gestures, allowing more screen content to be visible (Brasier et al., 2021). Qian et al. expanded on this and designed a free-hand manipulation system to manipulate AR objects using the rear-facing camera on a smartphone where they found through an iterative design process that haptic, visual, and sound cues assisted users in their AR tasks (Qian et al., 2019). Additionally, Loorak et al. presents hand and face interactions to interact with the on-screen UI, demonstrating users’ preference for this touchless interaction versus a touch-based interface when taking selfie photos (Loorak et al., 2019). The literature above illustrates the potential of touchless interactions, but they also share drawbacks they encountered when designing an AR experience for a single camera, as well as limited screen space with a narrow field of view.
2.2. Expanding the Interaction Space of Smartphones
@@
Expanding the camera view of smartphones has been explored to expand the user interaction space and environmental awareness. (Grubert et al., 2016) introduce GlassHands, which uses reflections from sunglasses to enable interaction around unmodified mobile devices. Yeo et al. present OmniSense, a smartphone system enhanced with an omnidirectional camera to provide better around-the-body awareness (Yeo et al., 2023). These approaches expand the field of view of mobile devices, allowing for more immersive, spatially demanding interactions – these findings share common insights with literature centered on expanding the input space of head-mounted displays with multi-camera configurations (Zhu and Grossman, 2020; Surale et al., 2019; Arora et al., 2018; Sakashita et al., 2023). Cam-2-Cam similarly seeks to broaden the interaction capabilities of smartphones, but focuses on using the existing front and rear cameras to capture different types of input and output without additional hardware.
2.3. Dual-Camera Applications in Mobile Tracking and Smartphone AR
@@
Prior work has demonstrated how smartphone cameras working synchronously can enhance real-world conditions experiences by providing additional sensing and improving environmental understanding. For instance, Babic et al. introduced SIMO, a system that uses the front camera for face and head tracking and the rear camera to estimate depth from a distant display to approximate body tracking (Babic et al., 2020). Nagai et al. presented HandyGaze, which uses gaze and head tracking via the front camera combined with the rear camera and pre-scanned 3D model of the environment to localize users’ gaze direction to identify artwork and present relevant information on a website in a mobile browser (Nagai et al., 2022).
@@
Zhao et al. used both the front-facing and rear-facing cameras to capture more accurate ambient lighting, enabling more realistic shaders and textures for AR try-on applications (Zhao et al., 2023). To the best of our knowledge, this is the only prior work that explicitly employs a dual-camera configuration for augmented reality on smartphones. However, while Zhao et al. focus on visual fidelity, the role of dual-camera setups for smartphone AR interaction techniques remains largely unexplored. Our work addresses this gap by investigating the interaction design potential of dual-camera configurations, extending the interaction space by delegating specific functions to each camera, without modifying the smartphone’s form factor.
@@
or shared input and content camera capture, and user values.
3. Cam-2-Cam Design
Face TriggAR, Mouth Craft, and Mirror ThrowAR serve as design probes for the Cam-2-Cam interaction concept. Each application expands the smartphone AR input space by delegating the front-facing feed gestural input and the rear camera feed to situate AR content. We delegated the rear-facing camera to display the resulting AR elements from users’ gestural actions captured by the front-facing camera because AR research and mainstream AR applications are primarily concerned with interacting with the physical world in front of the user (i.e., along the line of sight of the user). Additionally, our implemented applications incorporate audio, visual, and haptic feedback to enhance user engagement and provide immediate multimodal confirmation of actions. This design approach aligns with prior mobile AR research (Ma et al., 2023; Qian et al., 2019, 2020b; Wu and Li, 2024) emphasizing the benefits of multimodal feedback and touchless gestures in improving screen occlusion caused by limited screen real estate and enhancing user perceived control in smartphone-based AR applications.
3.1. Face Gesture Task Preferences
@@
Figure 2. Bar chart shows the number of participants selecting each gesture as their preferred dual-camera AR facial input after trying out projectile shooting and block object placement prototypes.
@@
3.2. Mobile Platform Considerations
@@
To enable hand gestures for free-hand manipulation of AR content we used the Portal-ble free-hand AR library due to its efficacy in manipulating AR objects as shown in several studies (Qian et al., 2019, 2020b; Ma et al., 2023). The Portal-ble library is only supported on Android, so we implemented Mirror ThrowAR application on Android (using the latest Google Pixel 9) rather than iOS.
@@
3.3. Face TriggAR
Face TriggAR is developed with Swift, using frameworks for AR mobile scenes (ARKit and RealityKit (Inc., 2024)), built for iOS smartphones and later deployed on the iPhone 15. It uses world, motion, and scene tracking with a single multi-camera capture session with simultaneous front and rear-cameras for a paintball target practice simulation.
3.3.1. Scene Recognition and Target Generation:
@@
Figure 3. Face TriggAR: (1) Displays a zoomed-in view of the user performing the wink gesture registered by the front-facing camera. (2) Shows the resulting projectiles (virtual paintballs) being shot along the user’s line of sight and visualized in the rear camera’s view.
3.3.2. Triggering Projectile Launches:
@@
3.4. Mouth Craft
Mouth Craft is an AR application designed for the iPhone 15 that uses ARKit and RealityKit frameworks with Swift to create a touchless interaction system driven by mouth gestures for object placement. Users can build virtual 3D structures by placing Minecraft-styled (Studios, 2024) grass blocks into their scene environment with the front-facing camera continuously tracking facial expressions. This design allows users to act in real-time without transition delays, making the AR environment feel more immediate and responsive. When a user opens their mouth, the blocks are situated in the AR scene and rendered to the rear camera’s feed.
3.4.1. Mouth Input:
@@
Figure 4. Mouth Craft: Starting from left to right, we demonstrate the (1) user performing the ‘mouth open’ gesture that is captured within the FOV of the front-facing camera and a detected plane is captured within the FOV of the rear-facing camera. (2) A block is placed on the detected plane and situated in the physical space captured by the rear-facing camera.
3.4.2. Block Scene Placement:
@@
3.5. Mirror ThrowAR
3.5.1. Throwing Implementation:
Mirror ThrowAR builds upon an open-source smartphone AR system, Portal-ble, to enable free-hand throwing detection (Qian et al., 2019). The system uses the Google Pixel 9’s front-facing and rear-facing camera to track the user’s spatial hand positions and convert them to 3D coordinates in the AR application in real time. These coordinates are used to detect overhand throwing gestures for our experiment.
We consider a three-dimensional Cartesian system, where $x$ is the horizontal axis, $y$ is the vertical axis relative to the virtual target generated on the wall, and $z$-axis is perpendicular to both $x$ and $y$, pointing outwards away from the cameras making the $z$-axis the forward axis in respect to the target location. Given these three axes, we can define three perpendicular planes: the $x$-$y$ plane, which describes the wall where the target is anchored; the $y$-$z$ plane that defines the vertical space between the participant and the target; and the $x$-$z$ plane, which identifies the horizontal space between the participant and the wall (see equation 1). We used ballistic motion calculations adapted from (Chudinov, 2014) :
(1) $\Theta_{i}=\arctan\frac{\left(S_{i}^{{}^{2}}+\sqrt{S_{i}^{4}-G\left(Gx^{2}+2S_{i}^{2}\right)}\right)}{G}$
@@ 3.5.1. Throwing Implementation:
Figure 5. Mirror ThrowAR users hold their hand out in a pinch or ball grabbing gesture to generate a Poké Ball fixed to their hand visible on their smartphone. Then, they throw the ball toward their smartphone into the rear camera AR scene to try to hit the Pikachu target.
4. Formative Evaluation
4.1. Participants
@@
Our recruitment goal was 30 participants as this is often sufficient to reach saturation of themes in qualitative experiments, and a higher number typically results in more nuanced insights (Guest et al., 2006; Iftikhar et al., 2023). The recruited participants included 12 females and 18 males between the ages of 18 and 60 years old ($\mu=26.97$, $\sigma=9.41$). Participants also provided their experience with smartphone AR during our semi-structured interview and rated on a 5-point Likert scale (1 = No Experience, 5 = Very Experienced). The resulting levels of experience can be summarized as ($\mu=1.93$, $\sigma=0.89$).
4.2. Study Protocol
@@
Each application session began with experimenters demonstrating the functions of the applications until participants felt comfortable trying on their own without training and starting the trial. In Face TriggAR, participants first scanned a QR code anchor image to generate a fixed AR target, spending the initial moments aiming at this anchored target. Afterwards, they could manually reset the target location such as on a table or coach, allowing them to explore aiming from various angles and even move behind the target to observe it from different perspectives. Using a wink gesture to aim and shoot at each target, participants had two minutes to interact with the scenes.
@@
4.3. Data Analysis
Experimenters conduct semi-structured interviews with the following qualitative questionnaire based on the AR Design Heuristics presented by Endsley et al. and tailored to fit our research questions displayed in table 1 (Endsley et al., 2017). The first and third authors carried out the thematic analysis by initially performing an open coding procedure followed by axial coding independently. Afterward, they reconciled differences in the final code groups and formed themes as seen in table 2.
Survey Questions AR Design Heuristics (Endsley et al., 2017)